# 1st Chenchen Wang College of Information Science

### v1

@@ IV-B Simulation of a 1D Parabolic PDE
When dealing with 1D Parabolic PDE, the implicit method is employed, which enhances accuracy and reduces computation. We employ the $dx=0.01$ over the time interval $[0,1]$ with a temporal step size of $\Delta t=0.001$. The coefficient $\lambda(x)$ is sampled from the Chebyshev polynomial $\lambda(x)=50\cos\big{(}\gamma\cos^{-1}(x)\big{)}$, where $\gamma\sim U[8,12]$. We generate a dataset comprising $6\times 10^{5}$ samples, which were produced by combining 100 different $\lambda(x)$ and 60 different $u_{0}(x)$. During the data collection process, data was collected every 100 time steps, its DeepONet training process is the same as that for the 1D Hyperbolic PDE. During testing, we extend the time $T$ to 3 seconds to observe the effects.
The DeepONet approximation results are shown in Fig. 8, illustrating the boundary control $U(t)$ and the $L_{2}$ norm of the state under the backstepping controller and the NO-based controller for $\gamma=9$. It can be seen that the state converges to zero under both controllers.
Figure 8: (a) Control input and (b) $L_{2}$ norm of the state $u(x,t)$ for hyperbolic PDE with $\gamma=9$ and $u_{0}(x)=9$ under the backstepping controller and the DeepONet-approximated controller.
The same $\gamma$ value is used, and the initial state $u_{0}(x)$ is randomly selected during the RL, the RL training process is the same as that for 1D Hyperbolic PDE. Fig. 4(b) shows the reward curves for SAC, NOSAC, and NOSAC_training on 1D Parabolic PDE, where NOSAC_training achieves faster reward growth.
We apply the NOSAC_training controller to the 1D parabolic PDE and compare its performance with the backstepping, SAC, and NOSAC, as shown in Fig. 9. The proposed NOSAC_training controller exhibits less overshoot. Fig. 10 plots the control effort and the $L_{2}$ norm of the state, showing that NOSAC_training is optimal in terms of overshoot and convergence speed, while the backstepping controller achieves a lower steady-state error.
@@
To evaluate robustness under model mismatch, we simulate a system with $\gamma=8.5$ using the three LR_based controllers trained at $\gamma=9$ and the backstepping controller designed for $\gamma=9$. The results in Fig. 11 demonstrate that NOSAC_training achieves superior robustness, outperforming the other controllers in overshoot, convergence rate, and steady-state error. Note that the backtepping controller performs worse for the parabolic PDE than the hyperbolic PDE under model mismatch.
@@ V Conclusion
This paper develops a novel neural-operator-embedded SAC architecture for PDE control, which incorporates the knowledge of rigorous control through a DeepONet that approximates the backstepping controller. We apply this learning-based control methodology to both a 1D hyperbolic and a 1D parabolic unstable PDE, and evaluate its performance in terms of overshoot, convergence rate, and steady-state error, compared to the backstepping controller, standard SAC, and SAC embedded with a non-pretrained DeepONet. Simulations on both systems demonstrate that the proposed method outperforms the other three controllers. The robustness against model mismatch is also evaluated by changing the system coefficient to a value different from that used for control design. The results show that the proposed method achieves superior performance compared to the other approaches, owing to the pre-trained DeepONet’s ability to adapt to varying coefficient functions and generate appropriate control signals in response to these changes. Future research will explore integrating safety control with this learning-based approach.

### v2

@@ IV-B Simulation of a 1D Parabolic PDE
When dealing with 1D Parabolic PDE, the implicit method is employed, which enhances accuracy and reduces computation. We employ the $dx=0.01$ over the time interval $[0,1]$ with a temporal step size of $\Delta t=0.001$. The coefficient $\lambda(x)$ is sampled from the Chebyshev polynomial $\lambda(x)=50\cos\big(\gamma\cos^{-1}(x)\big)$, where $\gamma\sim U[8,12]$. We generate a dataset comprising $6\times 10^{5}$ samples, which were produced by combining 100 different $\lambda(x)$ and 60 different $u_{0}(x)$. During the data collection process, data was collected every 100 time steps, its DeepONet training process is the same as that for the 1D Hyperbolic PDE. The total training process takes about 190 minutes.
The DeepONet approximation results are shown in Fig. 8, illustrating the boundary control and the $L_{2}$ norm of the state under the backstepping controller and the NO-based controller for $\gamma=9$. It can be seen that $U_{NO}$ remains almost indistinguishable from $U$, and the state converges to zero under both controllers($\|u_{NO}-u\|_{L_{2}}$ is less than $10^{-4}$ after about $0.5$ seconds).
Figure 8: (a) Control input and (b) $L_{2}$ norm of the state $u(x,t)$ for parabolic PDE with $\gamma=9$ and $u_{0}(x)=9$ under the backstepping controller and the DeepONet-approximated controller.
In the RL training, we set $\gamma=9$, while the other settings are consistent with those used in the hyperbolic PDE. The training processes of NOSAC_training, NOSAC, and SAC take approximately 20 minutes, 22 minutes, and 14 minutes, respectively. Fig. 4(b) shows the reward curves for the proposed method and other two baseline algorithms on 1D Parabolic PDE, demonstrating that NOSAC_training exhibits the fastest reward growth and convergence.
We apply the NOSAC_training controller to the 1D parabolic PDE and compare its performance with the backstepping, SAC, and NOSAC, as shown in Fig. 9. The proposed NOSAC_training controller exhibits less overshoot. Fig. 10 plots the control effort and the $L_{2}$ norm of the state, showing that NOSAC_training is optimal in terms of overshoot and convergence speed, conversely, the backstepping controller exhibits enhanced steady-state convergence.
@@
To evaluate robustness under model mismatch, we simulate a system with $\gamma=8.5$ using the three LR_based controllers trained at $\gamma=9$ and the backstepping controller designed for $\gamma=9$. The results in Fig. 11 demonstrate that NOSAC_training achieves superior robustness, outperforming the SAC and NOSAC controllers in overshoot, convergence speed, and steady-state error.
@@ V Conclusion
This paper proposes a novel neural-operator-embedded SAC architecture for PDE control, integrating rigorous control knowledge via a DeepONet approximating the backstepping controller. Applied to 1D hyperbolic and parabolic unstable PDEs, it is evaluated against backstepping control, standard SAC, and SAC with non-pretrained DeepONet in terms of overshoot, convergence rate, and steady-state error. Simulations show superior performance. Robustness to model mismatch (via modified system coefficients) is validated, with pre-trained DeepONet enabling adaptation to coefficient variations and generating appropriate control signals, ensuring better performance. Future work will explore integrating safety control with this learning-based approach.